OpenStudy (brucebaner):

What is the value of expression? 5^−8 • 5^−4 A.5^32 B.5^4 C.5^-4 D.5^-12

OpenStudy (brucebaner):

@kobeni-chan

OpenStudy (brucebaner):

@DanJS

OpenStudy (danjs):

recall a^x * a^y = a^x+y)

OpenStudy (brucebaner):

sure

OpenStudy (danjs):

for the same base to exponents, you can combine them by adding the exponents

OpenStudy (danjs):

when the like bases are multiplied

OpenStudy (brucebaner):

Its c

OpenStudy (brucebaner):

i did the problem and i got C

OpenStudy (danjs):

\[5^{-8} * 5^{-4} = 5^{-8 + (-4)}\]

OpenStudy (danjs):

not really concerned with the answer, more so of the process you did to get to the answer

OpenStudy (danjs):

so the answer is 5 to the -8+(-4) power

OpenStudy (brucebaner):

ohh its B

OpenStudy (danjs):

Just remember this \[A ^{x} * A ^{y} = A ^{x+y}\]

OpenStudy (danjs):

has to be the same base to apply that, in this case it was 5

OpenStudy (brucebaner):

so it b

OpenStudy (danjs):

-8+(-4) = -8-4 = -12

OpenStudy (brucebaner):

thx

OpenStudy (danjs):

5^(-12)